A particle of radius ' r' sliding inside a hemispherical cup of radius r. What os the expression for time period

Grade:12th pass

1 Answers

Kaushal Negi
4 years ago
The time period for revolution inside the hemispherical cup of radius r is given by the expression
T = 2\pi \sqrt{} \frac{l}{g}
l = 2pi r (circumference)
T = 2\pi \sqrt{} \frac{2\pi r}{g}